For his art class, Bryce bought 2 kg of dark clay and 3 kg of light clay, paying $22 for the clay. He later needed 1 kg of dark clay and 2 kg of light clay,costing $13 altogether. How much more per kg does the the dark clay cost than the light clay, in dollars


Answer by palanisamy(496)

Let the cost of dark clay = $ x per kg
and the cost of light clay = $ y per kg
Bryce bought 2 kg of dark clay and 3 kg of light clay, paying $22 for the clay
So we have, 2x+3y = 22 ...(1)
Then he bought 1 kg of dark clay and 2 kg of light clay,costing $13
x+2y = 13 ...(2)
(2)*2 => 2x+4y = 26 ...(3)
(1)-(3)=> 3y-4y = 22-26
-y = -4
y = 4
Substituting in (1), we get, 2x+3*4 = 22
2x = 22-12
2x = 10
x = 10/2 = 5
Therefore, the cost of dark clay = $ 5 per kg
and the cost of light clay = $ 4 per kg